OpenStudy (anonymous):

hmmmmmmmmmm... 1 + 1=?

OpenStudy (anonymous):

WINDOW!!!

OpenStudy (anonymous):

3

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

no =/ unfortunately...it's 2 =/

myininaya (myininaya):

no its 10

myininaya (myininaya):

binary

OpenStudy (anonymous):

..touche..

OpenStudy (anonymous):

i thought it is eleven. isn't it?

myininaya (myininaya):

10=2 so whatever 1+1=2=1^2^1+0*2^0=10